# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v1

@@ 5.1 Model and parameter specification
Let $n=2$ and suppose the energy spot price process ${\bf S}(t)$, $t\in[0,T_{o}]$ with terminal time $T=T_{o}$, is as specified in Section 4.1, where the LDOUP is ${\bf X}\sim{OU-WVAG}^{2}(\lambda,a,\alpha,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ with BDLP ${\bf Z}\sim{WVAG}^{2}(a,\alpha,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$.
@@
$\begin{gathered}\lambda=18.25,\quad a=3.15854,\quad\boldsymbol{\alpha}=\begin{pmatrix}\alpha_{1}\\ \alpha_{2}\end{pmatrix}=\begin{pmatrix}0.17183\\ 0.28494\end{pmatrix},\\ \boldsymbol{\mu}=\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix}=\begin{pmatrix}-0.03071\\ -0.20335\end{pmatrix},\quad{\boldsymbol{\Sigma}}=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{12}&\Sigma_{22}\end{pmatrix}=\begin{pmatrix}0.24598&0.19452\\ 0.19452&0.17045\end{pmatrix},\end{gathered}$ (5.5)
and $\boldsymbol{\eta}=-\boldsymbol{\mu}$. We refer to the seasonality and LDOUP parameters together as the model parameters. It is necessary to choose $\boldsymbol{\eta}$ such that $\mathbb{E}[{\bf Z}(1)]=\overline{\boldsymbol{\mu}}={\bf 0}$ so that the model is identifiable, otherwise both $\boldsymbol{\eta}$ and the intercept $(b_{10},b_{20})$ act as location parameters. Here, this choice is equivalent to $\boldsymbol{\eta}=-\boldsymbol{\mu}$. The initial value ${\bf X}_{0}$ is an independent random vector in the stationary distribution.
The market price of risk is
@@
The Carr-Madan dampening parameter for call options on $S_{1}$ and $S_{2}$ is $\epsilon=2.5$, which satisfies the conditions $(\epsilon+1){\bf e}_{1},(\epsilon+1){\bf e}_{2}\in{\cal D}_{{\bf Z}}$. The Hurd-Zhou dampening parameter is $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})=(-3.5,1)$, which satisfies the conditions $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$ and $-\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$.
@@
In theory, any dampening parameter satisfying the above-mentioned condition has no effect on the call and spread option prices. In practical terms, when numerically evaluating the pricing formulas, we check that changing to nearby dampening parameters has virtually no effect, and any choice satisfying this will do. Further discussions on the choice of dampening parameter is given in (AlSc2018, , Section 4.2).  $\Box$
@@ 5.2 Simulation
The observations of the LDOUP are ${\bf X}(t_{i})$ $t_{i}=i\Delta$, $i=0,1,\dots,m$, where $m=\lfloor{T_{e}/\Delta}\rfloor=2000$. These are simulated using the method in (Lu21, , Section 4.2) with $\widetilde{n}=1000$ subintervals. This means we use (2.11), where ${\bf Z}^{*}(\Delta)$ is simulated by discretizing the stochastic integral in (2.10) over $[0,\lambda\Delta]$ into $\widetilde{n}$ subintervals. Since the initial value ${\bf X}_{0}$ is assumed to be in its stationary distribution, which exists by (Lu21, , Proposition 1(i)), ${\bf X}_{0}$ is simulated by 20% burn-in.
@@ 5.3 Estimation
$L(\boldsymbol{\vartheta},{\bf x}_{0},\dots,{\bf x}_{m})=e^{mn\lambda\Delta}\prod_{k=1}^{m}f_{{\bf Z}^{*}(\Delta)}(e^{\lambda\Delta}{\bf x}_{k}-{\bf x}_{k-1}),$ (5.6)
$f_{{\bf Z}^{*}(\Delta)}$ is the pdf of ${\bf Z}^{*}(\Delta)$, computed by taking the Fourier inversion of the characteristic function $\Phi_{{\bf Z}^{*}(\Delta)}(\boldsymbol{\theta})=\exp(\kappa_{{\bf Z}^{*}(\Delta)}({\rm i}\boldsymbol{\theta}))$, and $\kappa_{{\bf Z}^{*}(\Delta)}$ is given by Corollary 4.6. The numerical Fourier inversion method is described in (MiSz17, , Section 4). Using the notation there and in (Lu21, , Section 5.1), we set the number of grid points along each axis to be $N=2^{12}$ for 1-dimensional FFT and $N=2^{8}$ for 2-dimensional FFT, and the spacing $h_{k}$ is chosen such that $\bar{x}_{k}=30\sqrt{\operatorname{Var}(Z^{*}_{k}(\Delta))}$, where $2\bar{x}_{k}$ is the width of the ${\bf x}$-grid along the $\bar{x}_{k}$ axis on which the pdf is outputted. The factor $f_{{\bf X}(0)}({\bf x}_{0})$ has not been included in (5.6) as a single observation in the likelihood is negligible in practice for large sample sizes $m$, however to be more correct, Lu21 does include it and shows how it can be computed.
@@ 5.4 Calibration
Next, we estimate the market price of risk ${\bf h}$, which is referred to as calibrating ${\bf h}$ or calibrating the model. This is done by solving
@@
Our approach estimates $\boldsymbol{\vartheta}$ under $\mathbb{P}$ and calibrates ${\bf h}$ under $\mathbb{Q}_{{\bf h}}$. This is in contrast to other methods, such as in Gardini and Sabino gs22 , where the mean-correcting martingale measure is used. This requires that the model be specified under a risk-neutral measure $\mathbb{Q}$ rather than $\mathbb{P}$, and then calibrated to forward and call prices. An issue then arises with the estimation of joint parameters, such as those affecting the correlation, which should be calibrated to the prices of other bivariate energy derivatives, but these may not be readily traded in the market, while estimating them using historical correlations under $\mathbb{P}$ is inconsistent with the model. In contrast, our model combines both the observed energy prices ${\bf S}$ under $\mathbb{P}$ and univariate energy derivative prices to fit the model in a theoretically sound way. In fact, it is unclear whether this model can be estimated only by calibrating to derivative prices, as it may be the case that $\boldsymbol{\vartheta}$ and ${\bf h}$ would become unidentifiable. This issue arises if energy prices follow the simpler Black-Scholes model.
Interestingly, $h_{2}$ affects the distribution of $Z_{1}$ under $\mathbb{Q}_{{\bf h}}$ as noted in Remark 4.4, so it seems possible in principle to calibrate ${\bf h}$ to energy derivatives on $S_{1}$ only, which is not possible in the model of Gardini and Sabino.
$\Box$
@@ 5.5 Pricing
Finally, we discuss the FFT method for numerically evaluating the spread option pricing formula in (4.16), which can be applied using the true or fitted model. While the implementation details are already given in HZ10 and also AlSc2018 , here we provide additional explanation on how we obtain the spread option price for a panel of strike prices $K_{i}$, with a single FFT evaluation.
Consider a 2-dimensional grid in the $\boldsymbol{\theta}$-domain and ${\bf x}$-domain. This is specified by $N$, the number of grid points along each axis, and $\overline{\theta}_{k}$, $k=1,2$, where $2\overline{\theta}_{k}$ is the width of the grid along the $\theta_{k}$-axis. If $\overline{\theta}_{1}=\overline{\theta}_{2}$, denote this value by $\overline{\theta}$. The grid are matrices, which we denote with a minor abuse of notation also as $\boldsymbol{\theta}\in\mathbb{R}^{{N}\times{N}}$ and ${\bf x}\in\mathbb{R}^{{N}\times{N}}$, which have entries
@@
where the $\theta_{k}$-spacing is $\delta_{\theta_{k}}=2\overline{\theta}_{k}/N$ and the $x_{k}$-spacing is $\delta_{x_{i}}=2\pi/(N\delta_{\theta_{i}})$ (note that $\overline{\theta}$, $\delta_{\theta}$, $\delta_{x}$ here are $\overline{u}$, $\eta$, $\eta^{*}$ in HZ10 ).
At the current time $t=T_{e}$, for the energy price ${\bf S}(t)=[{\bf x}]_{{\bf j}^{*}}$ on the ${\bf x}$-domain grid for some ${\bf j}^{*}$, it is shown in (HZ10, , Section 3) that, the price of the spread option is given by
$f_{t}({\bf S}(t),1)\approx\frac{e^{-rT}}{(2\pi)^{2}}\delta_{\theta_{1}}\delta_{\theta_{2}}N^{2}(-1)^{j^{*}_{1}+j^{*}_{2}}e^{-\langle\boldsymbol{\epsilon},[{\bf x}]_{{\bf j}^{*}}\rangle}[\mathscr{F}^{-1}\{{\bf H}\}]_{{\bf j}^{*}}$ (5.8)
@@
$[{\bf H}]_{\bf i}=(-1)^{j_{1}+j_{2}}\Phi([\boldsymbol{\theta}]_{\bf i}+{\rm i}\epsilon)\hat{P}([\boldsymbol{\theta}]_{\bf i}+{\rm i}\epsilon).$
@@
Now suppose we want to price the spread option for a general strike price $K>0$ using (4.17), then we require ${\bf S}(t)/K=[{\bf x}]_{\bf j}$ hold for some ${\bf j}$, that is, it is on the ${\bf x}$-grid. If ${\bf S}(t)/K$ is bounded by the ${\bf x}$-grid but not on it, taking the nearest grid point can lead to sizable pricing errors, so other approaches are considered.
@@
    Using bicubic interpolation of the ${\bf x}$-grid to compute to $f_{t}({\bf S}(t)/K,1)$.
@@
    Adjust the ${\bf x}$-grid so that ${\bf S}(t)/K$ is on it. Do this by setting ${\bf j}^{*}$ as the index of the grid point nearest ${\bf S}(t)/K$, and then adjust the values of $\delta_{\theta_{1}},\delta_{\theta_{2}}$ such that ${\bf S}(t)/K=[{\bf x}]_{{\bf j}^{*}}$ holds.
To compute the price of a spread option $f_{t}({\bf S}(t),K_{i})$ for a panel of strike prices $K_{i}$, $i=1,\dots,q$, on a single FFT evaluation, we do both. For a particular strike price of $K$, we adjust the ${\bf x}$-grid so that ${\bf S}(t)/K$ is on it. Then a single FFT evaluation outputs $(f_{t}([{\bf x}]_{\bf j},1))\in\mathbb{R}^{N\times N}$, from which we obtain $f_{t}({\bf S}(t)/K_{i},1)$ using bicubic interpolation. Finally, $f_{t}({\bf S}(t),K_{i})=K_{i}f_{t}({\bf S}(t)/K_{i},1)$ by (4.17). The grid is always adjusted so that $K=3.6$ is on it.
@@ 6.1 Spread option pricing under true parameters
Table 1: Spread option price over various strike prices $K_{i}$ using the FFT method with $N=2^{11}$, $\overline{\theta}=80$ and the Monte Carlo method with $N=10^{6}$ and $\widetilde{n}=10^{5}$.
@@
Average $\text{Average error}=\frac{1}{q}\sum_{i=1}^{q}|O_{i}-E_{i}|,$
where $O_{i}$ is the true price of the spread option with strike price $K_{i}$, and $E_{i}$ is the price of the spread option computed in another way. We take the true price to be from the FFT method with $N=2^{11}$, $\bar{u}=80$, and it is compared with using the discretization parameters $N=2^{k}$, $k=6,7,\dots 10$, $\overline{\theta}=40,80$. Figure 2 shows how this affects the average error of the FFT method.
@@ 6.2 Spread option parameter sensitivities
Figure 3 shows the effect of the LDOUP parameters on the spread option price with strike price $K=3.6$ and all others held constant. The parameters $a,\alpha_{1},\alpha_{2},\Sigma_{12}$ are plotting over their full range subject to the parameter constraints, the same is also true for the lower limit of $\Sigma_{11},\Sigma_{22}$.
@@ 6.3 Fitting the model
We consider a simulation study based on 1000 simulated sample paths. Recalling the details from Section 5, each sample path has 2001 observations for estimating the model parameters, and 20 forward prices or call prices for calibration.
@@
Figure 4: Sample path of the log price $Y_{k}(t)$ (black), true seasonality function $\Lambda_{k}(t)$ (green) and estimated seasonality function $\widehat{\Lambda}_{k}(t)$ (red) for $k=1,2$
@@
Table 2 gives the true value, mean estimate, estimated standard error, and estimated relative root mean square error (RRMSE) of the seasonal parameters ${\bf b}$, LDOUP parameters $\boldsymbol{\vartheta}$, market prices of risk ${\bf h}^{f}=(h_{1}^{f},h_{2}^{f})$ and ${\bf h}^{c}=(h_{1}^{c},h_{2}^{c})$ from calibrating to forward and call prices, respectively, and their corresponding risk-neutral drifts $\widetilde{\boldsymbol{\mu}}^{f}=(\widetilde{\mu}_{1}^{f},\widetilde{\mu}_{2}^{f})$, $\widetilde{\boldsymbol{\mu}}^{c}=(\widetilde{\mu}_{1}^{c},\widetilde{\mu}_{2}^{c})$
@@
Table 2: True value, mean estimate, standard error, and RRMSE for the seasonal parameter, LDOUP parameters, market prices of risk, risk-neutral drifts estimated over 1000 simulation with 2001 observations and 20 calibration instruments per simulation.
From these results, most model parameters are estimated with moderately low RRMSE of approximately 10–30% other than those with true values near 0. Related to this, note that errors in estimating the model parameters are compensated for by the calibration. For instance, $h_{1}^{c}$ and $h_{2}^{c}$ have RRMSEs of 589% and 2612%, which can take them far from their true value, so that $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ can have small RRMSEs of 49% and 12%, close to their true value. In fact, on an absolute scale as seen by the vertical line in Figure 5, the differences between the true and estimated values of $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ are negligible. Ultimately, it is $\widetilde{\boldsymbol{\mu}}$ that has a first-order effect on the energy derivative prices, and the fitted models should primarily be judged on how accurately it prices spread options.
Figure 6: The true price of call options on $S_{1}$ (black), and based on estimated model parameters, the call price using the true market price of risk ${\bf h}$ (green), using the estimated market price of risk $\widehat{\bf h}^{c}$.
Figure 6 shows the results of calibrating to call prices. Based on the estimated model parameters, the estimated value $\widehat{\bf h}^{c}$ yields a very close fit to the true prices of call options on $S_{1}$, but the true value ${\bf h}$ does not. While Figures 5 and 6 are based on a single simulation, these are typical results seen in most. Only results for calibrating to call prices on $S_{1}$ are plotted, although we calibrated to call prices on $S_{1}$ and $S_{2}$. A similarly close match holds for call prices on $S_{2}$, while calibration to forwards does not produce such a close match.

### v2

@@ 5.1 Model and parameter specification
Let $n=2$ and suppose the energy spot price process ${\bf S}(t)$, $t\in[0,T_{o}]$ with terminal time $T=T_{o}$, is as specified in Section 4.1, where the LDOUP is ${\bf X}\sim{OU\text{-}WVAG}^{2}(\lambda,a,\alpha,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ with BDLP ${\bf Z}\sim{WVAG}^{2}(a,\alpha,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$.
@@
[…]
and $\boldsymbol{\eta}=-\boldsymbol{\mu}$. We refer to the seasonality and LDOUP parameters together as the model parameters. It is necessary to choose $\boldsymbol{\eta}$ such that $\mathbb{E}[{\bf Z}(1)]=\overline{\boldsymbol{\mu}}={\bf 0}$, which is equivalent to $\boldsymbol{\eta}=-\boldsymbol{\mu}$, so that the model is identifiable, otherwise both $\boldsymbol{\eta}$ and the intercept $(b_{10},b_{20})$ act as location parameters. The initial value ${\bf X}_{0}$ is an independent random vector in the stationary distribution.
The MPR is
@@
The Carr-Madan damping parameter for call options on $S_{1}$ and $S_{2}$ is $\epsilon=2.5$, which satisfies the conditions $(\epsilon+1){\bf e}_{1},(\epsilon+1){\bf e}_{2}\in{\cal D}_{{\bf Z}}$. The Hurd-Zhou damping parameter is $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})=(-3.5,1)$, which satisfies the conditions $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$ and $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$.
@@
In theory, any damping parameter satisfying the above-mentioned condition has no effect on the call and spread option prices. In practical terms, when numerically evaluating the pricing formulas, we check that changing to nearby damping parameters has virtually no effect, and any choice satisfying this will do. Further discussions on the choice of damping parameter is given in (AlSc2018, , Section 4.2).  $\Box$
@@ 5.2 Simulation
The observations of the LDOUP ${\bf X}(t_{i})$ $t_{i}=i\Delta$, $i=0,1,\dots,m$, where $m=\lfloor{T_{e}/\Delta}\rfloor=2000$ are simulated with $\widetilde{n}=1000$ subintervals using the method in (Lu21, , Section 4.2). This means we use (2.11), where ${\bf Z}^{*}(\Delta)$ is simulated by discretizing the stochastic integral in (2.10) over $0,\lambda\Delta]$ into $\widetilde{n}$ subintervals. Since the initial value ${\bf X}_{0}$ is assumed to be in its stationary distribution, which exists by (Lu21, , Proposition 1(i)), ${\bf X}_{0}$ is simulated by 20% burn-in.
@@ 5.3 Estimation
$L(\boldsymbol{\vartheta},{\bf x}_{0},\dots,{\bf x}_{m})=e^{mn\lambda\Delta}\prod_{k=1}^{m}f_{{\bf Z}^{*}(\Delta)}(e^{\lambda\Delta}{\bf x}_{k}-{\bf x}_{k-1}).$ (5.6)
Here, $f_{{\bf Z}^{*}(\Delta)}$ is the pdf of ${\bf Z}^{*}(\Delta)$, which is computed by taking the Fourier inversion of the characteristic function $\Phi_{{\bf Z}^{*}(\Delta)}(\boldsymbol{\theta})=\exp(\kappa_{{\bf Z}^{*}(\Delta)}({\rm i}\boldsymbol{\theta}))$ using a version of Corollary 4.6 that is under $\mathbb{P}$, specifically,
$\kappa_{{\bf Z}^{*}(\Delta)}(\boldsymbol{\theta})={\rm i}\langle(e^{\lambda\Delta}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\kappa_{{\bf V}_{0}^{*}(\Delta)}(\boldsymbol{\theta})+\sum_{i=1}^{n}\kappa_{V_{k}^{*}(\Delta)}(\theta_{k}),\quad\boldsymbol{\theta}\in{\cal D}_{{\bf Z}},$
where the law of ${\bf V}_{0}$, $V_{k}$, $k=1,\dots,n$, under $\mathbb{P}$ are given in (2.5). The numerical Fourier inversion method is described in (MiSz17, , Section 4). Using the notation there and in (Lu21, , Section 5.1), we set the number of grid points along each axis to be $N=2^{12}$ for 1-dimensional FFT and $N=2^{8}$ for 2-dimensional FFT, and the spacing $h_{k}$ is chosen such that $\bar{x}_{k}=30\sqrt{\operatorname{Var}(Z^{*}_{k}(\Delta))}$, where $2\bar{x}_{k}$ is the width of the $\bf x$}-grid along the $\bar{x}_{k}$ axis on which the pdf is outputted. The factor $f_{{\bf X}(0)}({\bf x}_{0})$ has not been included in (5.6) as a single observation in the likelihood is negligible in practice for large sample sizes $m$, however to be more correct, Lu21 does include it and shows how it can be computed.
@@ 5.4 Calibration
Next, we estimate the MPR $\bf h$} by calibrating the model, which specifically solves
@@
Our approach estimates $\boldsymbol{\vartheta}$ under $\mathbb{P}$ and calibrates $\bf h$} under $\mathbb{Q}_{{\bf h}}$, in contrast to the approach using the mean-correcting martingale measure, such as in Gardini and Sabino gs22 . The latter requires the model be specified under a risk-neutral measure $\mathbb{Q}$ rather than $\mathbb{P}$, and then calibrated to forward and call prices. An issue then arises with the estimation of joint parameters, such as those affecting the correlation, which should be calibrated to the prices of other bivariate energy derivatives, but these may not be readily traded in the market, while estimating them using historical correlations under $\mathbb{P}$ is inconsistent with the model. In contrast, our method combines both the observed energy prices $\bf S$} under $\mathbb{P}$ and univariate energy derivative prices to fit the model in a theoretically sound way. In fact, it is unclear whether our model can be estimated only by calibrating to derivative prices, as it may be the case that $\boldsymbol{\vartheta}$ and $\bf h$} would become unidentifiable. This issue arises if energy prices follow the simpler Black-Scholes model.
Interestingly, $h_{2}$ affects the distribution of $Z_{1}$ under $\mathbb{Q}_{{\bf h}}$ as noted in Remark 4.4, so it seems possible in principle to calibrate $\bf h$} to energy derivatives on $S_{1}$ only, which is not possible in the model of Gardini and Sabino.   $\Box$
@@ 5.5 Pricing
Finally, we discuss the FFT method for numerically evaluating the spread option pricing formula in (4.17), which can be applied using the true or fitted model. While the implementation details are already given in HZ10 , and also AlSc2018 , here we provide additional explanation on how we obtain the spread option price for a panel of strike prices $K_{i}$, with a single FFT evaluation.
Consider a 2-dimensional grid in the $\boldsymbol{\theta}$-domain and $\bf x$}-domain. This is specified by $N$, the number of grid points along each axis, and $\overline{\theta}_{k}$, $k=1,2$, where $2\overline{\theta}_{k}$ is the width of the grid along the $\theta_{k}$-axis. If $\overline{\theta}_{1}=\overline{\theta}_{2}$, denote this value by $\overline{\theta}$. Then the $\boldsymbol{\theta}$-grid and $\bf x$}-grid are arrays whose entries are
@@
where the $\theta_{k}$-spacing is $\delta_{\theta_{k}}=2\overline{\theta}_{k}/N$ and the $x_{k}$-spacing is $\delta_{x_{k}}=2\pi/(N\delta_{\theta_{k}})$ (note that $\overline{\theta}$, $\delta_{\theta}$, $\delta_{x}$ here are $\overline{u}$, $\eta$, $\eta^{*}$ in HZ10 ).
At the current time $t=T_{e}$, for the log energy price $\log{\bf S}(t)=[{\bf x}]_{{\bf j}^{*}}$ on the $\bf x$}-grid for some $\bf j$$^{*}$, it is shown in (HZ10, , Section 3) that the price of the spread option is approximated by
$f_{t}({\bf S}(t),1)\approx\frac{e^{-rT}}{(2\pi)^{2}}\delta_{\theta_{1}}\delta_{\theta_{2}}N^{2}(-1)^{j^{*}_{1}+j^{*}_{2}}e^{-\langle\boldsymbol{\epsilon},[{\bf x}]_{{\bf j}^{*}}\rangle}[\mathscr{F}^{-1}\{{\bf H}\}]_{{\bf j}^{*}},$ (5.8)
@@
$[{\bf H}]_{\bf j}=(-1)^{j_{1}+j_{2}}\exp(\kappa^{\mathbb{Q}_{{\bf h}}}_{{\bf R}(t,T)}(-\boldsymbol{\epsilon}+{\rm i}[\boldsymbol{\theta}]_{\bf j}))\hat{P}([\boldsymbol{\theta}]_{\bf j}+{\rm i}\boldsymbol{\epsilon}).$
@@
Now suppose we want to price the spread option for a general strike price $K>0$ using (4.18), then we require $\log({\bf S}(t)/K)=[{\bf x}]_{\bf j}$ hold for some $\bf j$}, that is, it is on the $\bf x$}-grid. If $\log({\bf S}(t)/K)$ is bounded by the $\bf x$}-grid but not on it, taking the nearest grid point can lead to sizable pricing errors, so other approaches are considered.
@@
    Using bicubic interpolation on the $\bf x$}-grid to compute to $f_{t}({\bf S}(t)/K,1)$.
@@
    Adjust the $\bf x$}-grid so that $\log({\bf S}(t)/K)$ is on it. Do this by setting $\bf j$$^{*}$ as the index of the grid point nearest $\log({\bf S}(t)/K)$, and then adjust the values of $\delta_{\theta_{1}},\delta_{\theta_{2}}$ such that $\log({\bf S}(t)/K)=[{\bf x}]_{{\bf j}^{*}}$ holds.
To compute the price of a spread option $f_{t}({\bf S}(t),K_{i})$ for a panel of strike prices $K_{i}$, $i=1,\dots,q$, on a single FFT evaluation, we do both. For a particular strike price of $K$, we adjust the $\bf x$}-grid so that $\log({\bf S}(t)/K)$ is on it. Then a single FFT evaluation outputs $(f_{t}([{\bf x}]_{\bf j},1))\in\mathbb{R}^{N\times N}$, from which we obtain $f_{t}({\bf S}(t)/K_{i},1)$ using bicubic interpolation. Finally, $f_{t}({\bf S}(t),K_{i})=K_{i}f_{t}({\bf S}(t)/K_{i},\allowbreak 1)$ by (4.18). The grid is always adjusted so that $K=3.6$ is on it.
@@ 6.1 Spread option pricing under true parameters
Table 1: Spread option prices over various strike prices $K_{i}$ using the FFT method with $N=2^{11}$, $\overline{\theta}=80$ and the Monte Carlo method with $N=10^{6}$ and $\widetilde{n}=10^{5}$.
@@
$\text{Average error}=\frac{1}{q}\sum_{i=1}^{q}|O_{i}-E_{i}|,$
where $O_{i}$ is the true price of the spread option with strike price $K_{i}$, and $E_{i}$ is the price of the spread option computed in another way. We take the true price to be from the FFT method with $N=2^{11}$, $\bar{u}=80$, and it is compared with the prices using the discretization parameters $N=2^{k}$, $k=6,7,\dots 10$, $\overline{\theta}=40,80$. Figure 2 shows how this affects the average error of the FFT method.
@@ 6.2 Spread option parameter sensitivities
Figure 3 shows the effect of the LDOUP parameters on the spread option price with strike price $K=3.6$ and all others held constant. The parameters $a,\alpha_{1},\alpha_{2},\Sigma_{12}$ are plotted over their full range subject to the parameter constraints, the same is also true for the lower limit of $\Sigma_{11},\Sigma_{22}$.
@@ 6.3 Fitting the model
We consider a simulation study based on 1000 simulated sample paths. Recalling the details from Section 5, each sample path has 2001 observations for estimating the model parameters, and 40 forward prices or call prices for calibration.
@@
Figure 4: Sample path of the log price $Y_{k}(t)$ (black), true seasonality function $\Lambda_{k}(t)$ (green) and estimated seasonality function $\widehat{\Lambda}_{k}(t)$ (red) for $k=1,2$.
@@
Table 2 gives the true value, mean estimate, estimated standard error, and estimated relative root mean square error (RRMSE) of the seasonal parameters $\bf b$}, LDOUP parameters $\boldsymbol{\vartheta}$, MPRs ${\bf h}^{f}=(h_{1}^{f},h_{2}^{f})$ and ${\bf h}^{c}=(h_{1}^{c},h_{2}^{c})$ from calibrating to forward and call prices, respectively, and their corresponding risk-neutral drifts $\widetilde{\boldsymbol{\mu}}^{f}=(\widetilde{\mu}_{1}^{f},\widetilde{\mu}_{2}^{f})$, $\widetilde{\boldsymbol{\mu}}^{c}=(\widetilde{\mu}_{1}^{c},\widetilde{\mu}_{2}^{c})$
@@
Table 2: True value, mean estimate, standard error, and RRMSE for the seasonal parameters, LDOUP parameters, MPRs, risk-neutral drifts estimated over 1000 simulation with 2001 observations and 40 calibration instruments per simulation.
From these results, most model parameters are estimated with moderately low RRMSE of approximately 10–30%, other than those with true values near 0. Related to this, note that errors in estimating the model parameters are compensated for by the calibration. For instance, $h_{1}^{c}$ and $h_{2}^{c}$ have RRMSEs of 589% and 2612%, which can take them far from their true value, so that $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ can have small RRMSEs of 49% and 12%, close to their true value. In fact, on an absolute scale as seen by the vertical line in Figure 5, the differences between the true and estimated values of $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ are negligible. Ultimately, it is $\widetilde{\boldsymbol{\mu}}$ that has a first-order effect on the energy derivative prices, and the fitted models should primarily be judged on how accurately they prices spread options.
Figure 6: The true call prices on $S_{1}$ (black) over various strike prices $K_{i}$, and the predicted call price using the true MPR $\bf h$} (green) and calibrated MPR $\widehat{\bf h}^{c}$ (red).
Figure 6 shows the results of calibrating to call prices. Based on the estimated model parameters, the calibrated MPR $\widehat{\bf h}^{c}$ gives predicted call prices with a very close fit to the true call prices on $S_{1}$, but the true MPR $\bf h$} does not. While Figures 5 and 6 are based on a single simulation, these are typical results seen in most simulations. Only results for calibrating to call prices on $S_{1}$ are plotted, although we calibrated to call prices on $S_{1}$ and $S_{2}$. A similarly close match holds for call prices on $S_{2}$, while calibration to forwards does not produce such a close match.